The sum of the ages of the First year Math. teacher is 50 19 years ago from now.Ms. Santos' age will be twice the age of Ms. Valdez present age. How old are they now?


Answer by Earlsdon(6103)

We'll let S be Ms. Santos' present age and V be Ms. Valdez's present age.
At present:
1) S + V = 50 The sum of their ages is 50.
19 years from now:
2) S + 19 = 2V In 19 years, Ms. Santos will be twice Ms. Valdez present age.
Rewrite equation 1) as:
V = 50 - S Multiply both sides by 2 because you want to get 2V.
2V = 2(50 - S) = 100 - 2S Now substitute this for the 2V in equation 2).
2) S + 19 = 100 - 2S Solve this for S. Add 2S to both sides.
3S + 19 = 100 Subtract 19 from both sides.
3S = 81 Finally, divide both sides by 3.
S = 27
Ms. Santos' present age is 27 years
V = 50 - S
V = 50 - 27
v = 23
Ms. Valdez's present age is 23 years.